# Geometric Multi-color Message Passing Graph Neural Networks for Blood-brain Barrier Permeability Prediction

### v2

@@ 4 Results and Discussion
In this section, we present the experimental results of GMC-MPNN on the datasets $BBBP_{\text{cls}}^{\text{MolNet}}$, $BBBP_{\text{cls}}^{\text{B3DB}}$, and $BBBP_{reg}^{\text{B3DB}}$ coupled with various state-of-the-art models to evaluate its performance in BBBP.
4.1 Model Parametrization and Hyperparameter Optimization
Hyperparameter Search Space
Activation [RELU, LEAKYRELU, PRELU, TANH, SELU, ELU]
Aggregation [MeanAggregation, SumAggregation, NormAggregation]
Aggregation Norm Uniform distribution: [1, 200] (step size = 1)
Batch Size Discrete choices: [16, 32, 64, 128, 256]
Depth Random integer: [2, 6] (step size = 1)
Dropout [0.0, 0.05, 0.1, 0.15, $\dots$, 0.45]
FFN Hidden Dim Random integer: [300, 2400] (step size = 100)
FFN Num Layers Random integer: [1, 3]
Final LR Ratio Log-uniform distribution: [1e-2, 1]
Message Hidden Dim Random integer: [300, 2400] (step size = 100)
Init LR Ratio Log-uniform distribution: [1e-2, 1]
Max LR Log-uniform distribution: [1e-4, 1e-2]
Kernel type Exponential, Lorentz
Scaling factor ($\tau$) [0.5, 1.0, …, 10]
Characteristic distance ($\eta$) [0.5, 1.0, …, 20]
Table 2: Hyperparameter search space for GMC-MPNN.
We systematically optimize hyperparameters using a grid search strategy over a well-defined search space defined in Table 2. The search focuses on parameters such as activation functions, aggregation methods, model depth, dropout rates, feedforward network dimensions, learning rate schedules, and batch sizes.
Additionally, we also optimize WCS’s parameters to achieve robust and reliable atom-level descriptors. The WCS features are parametrized by $\beta$, $\kappa$, and $\tau$. Whereas $\beta$ specifies the kernel type, where we use the generalized exponential kernel ($\beta=E$) and the generalized Lorentz kernel ($\beta=L$), $\tau$ acts as a scaling factor that determines the characteristic distance $\eta_{kk^{\prime}}$ between atom types $k$ and $k^{\prime}$ using $\eta_{kk^{\prime}}=\tau(r_{k}+r_{k^{\prime}})$, where $r_{k}$ and $r_{k^{\prime}}$ are the van der Waals radii, and $\kappa$ is the power parameter of the kernel, which helps approximate the ideal low-pass filter (ILF) [30].
We then generate 1600 atom features, each corresponding to a unique combination of kernel parameters ($\beta$, $\kappa$, and $\tau$) selected from the grid search space $\tau\in[0.5,10]$ and $\kappa\in[0.5,20]$, incremented by 0.5. For each feature of GMC-MPNN, we evaluate the performance of the models trained with the 1600 feature sets, selecting the kernel configuration that yields the best performance on the validation set. This configuration is then used to benchmark the model on the test set. This Table 3 shows our best hyperparameters for the three datasets
Hyperparameter $BBBP_{\text{cls}}^{\text{MolNet}}$ $BBBP_{\text{cls}}^{\text{B3DB}}$ $BBBP_{\text{reg}}^{\text{B3DB}}$
Activation LEAKYRELU RELU PRELU
Aggregation SumAggregation MeanAggregation MeanAggregation
Aggregation Norm 57 4 41
Batch Size 32 32 16
Depth 5 3 3
Dropout 0.0 0.0 0.0
FFN Hidden Dim 900 700 2200
FFN Num Layers 2 2 1
Final LR Ratio 0.000544 9.42e-05 2.11e-05
Message Hidden Dim 2200 1300 1400
Init LR Ratio 5.70e-05 1.56e-04 6.37e-04
Max LR 0.00521 0.00385 6.38e-04
Kernel Type Exponential Exponential Lorentz
Scaling Factor ($\tau$) 0.5 5 7.5
Characteristic Distance ($\eta$) 17 12 16
Table 3: Best hyperparameter configurations for the BBBP classification datasets ($BBBP_{\text{cls}}^{\text{MolNet}}$, $BBBP_{\text{cls}}^{\text{B3DB}}$) and B3DB regression dataset ($BBBP_{\text{reg}}^{\text{B3DB}}$).
@@ 4.2 Classification Task Experiments
Figure 3: Performance comparison of different models on the $BBBP_{\text{cls}}^{\text{MolNet}}$ dataset. AUC scores for baseline methods are taken from Zhao et al. [38]. The red bars highlight our GMC-MPNN model.
@@
On the $BBBP_{\text{cls}}^{\text{MolNet}}$ dataset, as illustrated in Figure 3 and detailed in Table 4, GMC-MPNN achieves the best AUC-ROC score of 0.9704±0.0310.9704$\pm$ 0.0310.9704 ± 0.031, using kernel parameters $\beta=L$, $\kappa=17$, and $\tau=0.5$. This strong performance demonstrates the value of incorporating spatially-informed atomic features and long-range geometric interactions via weighted colored subgraphs. GSL-MPP (0.957±0.0080.957$\pm$ 0.0080.957 ± 0.008) and CD-MVGNN (0.933±0.0060.933$\pm$ 0.0060.933 ± 0.006) follow as the next-best models, while traditional message-passing models like MPNN (0.913±0.0410.913$\pm$ 0.0410.913 ± 0.041) and GraphConv (0.877±0.0360.877$\pm$ 0.0360.877 ± 0.036) perform less competitively.
@@
We next evaluate all models on the $BBBP_{\text{cls}}^{\text{B3DB}}$ dataset, which is considered more challenging due to its greater structural diversity and broader distribution of molecular scaffolds. As shown in Figure 4 and detailed in Table 5, GMC-MPNN again achieves the best performance with an AUC-ROC of 0.9685±0.01270.9685$\pm$ 0.01270.9685 ± 0.0127, confirming its robustness in generalizing across these more varied structures. Uni-Mol also performs strongly, achieving an AUC-ROC of 0.9452±0.01100.9452$\pm$ 0.01100.9452 ± 0.0110. Among conventional GNNs, D-MPNN achieves 0.8878±0.03810.8878$\pm$ 0.03810.8878 ± 0.0381, followed by GAT and GCN. Models like MPNN and Weave perform less favorably, with AUCs under 0.77, highlighting their limitations in capturing complex structural patterns under challenging generalization scenarios.
Table 5: Performance of various models on $BBBP_{\text{cls}}^{\text{B3DB}}$ dataset
@@
Overall, our GMC-MPNN model consistently outperforms all baselines on both benchmarks, achieving AUC-ROC scores of 0.9704±0.0310.9704$\pm$ 0.0310.9704 ± 0.031 on the MoleculeNet version and 0.9685±0.01270.9685$\pm$ 0.01270.9685 ± 0.0127 on the B3DB variant. The superior and consistent performance of GMC-MPNN across both benchmarks confirms the utility of incorporating geometry-aware atomic and subgraph features. These results underscore the importance of evaluating molecular property prediction models on diverse and carefully curated datasets to ensure reliable generalization in drug discovery pipelines.
@@ 4.3 Regression Task Experiments
Figure 5: Performance comparison of different models on the $\text{BBBP}_{\text{reg}}^{\text{B3DB}}$ dataset with RMSE and Pearson Correlation scores. The red bars showcase the performances of our geometric graph learning-based model, GMC-MPNN. We compare against several baseline and state-of-the-art models: GCN [39], NF [40], D-MPNN [32], Uni-mol [41], GIN [42], GAT [43], Weave [44], AttentiveFP [45], and MPNN [23].
Table 6: Performance of various models on $BBBP_{reg}^{\text{B3DB}}$ dataset. PCC stands for Pearson’s correlation coefficient.
@@
GMC-MPNN 0.4609$\pm$0.0651 0.7759$\pm$0.0699
GCN 0.6240$\pm$0.0252 0.6415$\pm$0.0498
NF 0.6325$\pm$0.0412 0.6145$\pm$0.0598
D-MPNN 0.6696$\pm$0.0536 0.6083$\pm$0.0504
GIN 0.6773$\pm$0.0604 0.5581$\pm$0.0651
Uni-mol 0.6747$\pm$0.0290 0.5354$\pm$0.0559
GAT 0.7011$\pm$0.0321 0.4538$\pm$0.0599
Weave 0.7334$\pm$0.0330 0.4339$\pm$0.0609
AttentiveFP 0.7444$\pm$0.0329 0.2992$\pm$0.0816
MPNN 0.7850$\pm$0.0387 -0.0400$\pm$0.1133
We evaluate the performance of our proposed GMC-MPNN model on the regression task using the $BBBP_{\text{reg}}^{\text{B3DB}}$ dataset. GMC-MPNN, which incorporates geometry-aware WCS descriptors and max-pooling, achieves the best results among all evaluated models. It records the lowest RMSE of 0.4609±0.06510.4609$\pm$ 0.06510.4609 ± 0.0651 and the highest Pearson’s correlation coefficient (PCC) of 0.7759±0.06990.7759$\pm$ 0.06990.7759 ± 0.0699. These results were obtained using kernel parameters $\beta=L$, $\kappa=16$, and $\tau=7.5$. The strong performance demonstrates the effectiveness of WCS features in capturing both spatial and chemical contexts critical for BBB permeability prediction.
Table 6 and Figure 5 present a comparative analysis across baseline models. Traditional graph-based models like GCN [39] and NF [40] follow with RMSEs of 0.6240±0.02520.6240$\pm$ 0.02520.6240 ± 0.0252 and 0.6325±0.04120.6325$\pm$ 0.04120.6325 ± 0.0412, and PCs of 0.6415±0.04980.6415$\pm$ 0.04980.6415 ± 0.0498 and 0.6145±0.05980.6145$\pm$ 0.05980.6145 ± 0.0598, respectively. D-MPNN [32] also shows competitive performance (RMSE 0.6696±0.05360.6696$\pm$ 0.05360.6696 ± 0.0536, PCC 0.6083±0.05040.6083$\pm$ 0.05040.6083 ± 0.0504).
Moderate results are observed from models such as GIN [42] (RMSE 0.67730.67730.6773, PCC 0.55810.55810.5581), Uni-Mol [41] (RMSE 0.67470.67470.6747, PCC 0.53540.53540.5354), and GAT [45] (RMSE 0.70110.70110.7011, PCC 0.45380.45380.4538). In contrast, Weave [44] and AttentiveFP [46] yield weaker outcomes, particularly AttentiveFP with a low PCC of 0.2992±0.08160.2992$\pm$ 0.08160.2992 ± 0.0816. MPNN [23] performs the worst, with the highest RMSE (0.7850±0.03870.7850$\pm$ 0.03870.7850 ± 0.0387) and a negative PCC (0.0400±0.1133-0.0400$\pm$ 0.1133- $0$ 0.1133), indicating poor generalization.
@@ 6 Conclusion
Ethics Statement
This study involved only publicly available datasets that contain chemical compound information and associated blood-brain barrier permeability measurements. No experiments involving human participants or animals were conducted. The datasets used (BBBPMolNet and B3DB) were previously published and curated; all data handling, model development, and analyses were performed under ethical research standards in computational biology and cheminformatics. No personally identifiable information or sensitive health data was used. The research was conducted with a focus on improving drug discovery pipelines for the public good, with no conflict of interest or involvement in unethical practices.

### v4

@@ 4 Results and Discussion
In this section, we present the experimental results of GMC-MPNN on the datasets $BBP_{\text{cls}}^{\text{MolNet}}, BB$$BBP_{\text{cls}}^{\text{B3DB}}$, and $BBBP_{reg}^{\text{B3DB}}$ coupled with various state-of-the-art models to evaluate its performance in BBBP.
@@ 4.1 Classification Task Experiments
Figure 3: Performance comparison of different models on the $BBBP_{\text{cls}}^{\text{MolNet}}$ dataset. The red bars highlight our GMC-MPNN model.
@@
On the $BBBP_{\text{cls}}^{\text{MolNet}}$ dataset, as illustrated in Figure 3 and detailed in Table 2, GMC-MPNN achieves the best AUC-ROC score of $0.947\pm 0.011$. This strong performance demonstrates the value of incorporating spatially-informed atomic features and long-range geometric interactions via weighted colored subgraphs. GSL-MPP [gslmpp] ($0.935\pm 0.039$) and CD-MVGNN [cdmvgnn] ($0.933\pm 0.006$) follow as the next-best models, while traditional message-passing models like MPNN [gilmer2017neural] ($0.913\pm 0.041$) and GraphConv [graphconv] ($0.877\pm 0.036$) perform less competitively.
@@
We next evaluate all models on the $BBBP_{\text{cls}}^{\text{B3DB}}$ dataset, which is considered more challenging due to its greater structural diversity and broader distribution of molecular scaffolds. As shown in Figure 4 and detailed in Table 3, GMC-MPNN again achieves the best performance with an AUC-ROC of $0.9212\pm 0.0261$, confirming its robustness in generalizing across these more varied structures. Following the top performers, a cluster of recent models demonstrates comparable results, including CD-MVGNN [cdmvgnn] ($0.8958\pm 0.0669$), CoMPT [compt] ($0.8896\pm 0.0039$), D-MPNN [heid2023chemprop] ($0.8878\pm 0.0381$), and GSL-MPP [gslmpp] ($0.8840\pm 0.0126$). These are followed by GAT [velickovic2017graph] and GCN [kipf2016semi]. Models like MPNN [gilmer2017neural] and Weave perform less favorably, with AUCs under 0.77, highlighting their limitations in capturing complex structural patterns under challenging generalization scenarios.
Table 3: Performance of various models on the $BBBP_{\text{cls}}^{\text{B3DB}}$ dataset, sorted by performance.
@@
Overall, our GMC-MPNN model consistently outperforms all baselines on both benchmarks, achieving AUC-ROC scores of $0.947\pm 0.011$ on the MoleculeNet version and $0.9212\pm 0.0261$ on the B3DB variant. The superior and consistent performance of GMC-MPNN across both benchmarks confirms the utility of incorporating geometry-aware atomic and subgraph features. These results underscore the importance of evaluating molecular property prediction models on diverse and carefully curated datasets to ensure reliable generalization in drug discovery pipelines.
@@ 4.2 Regression Task Experiments
Figure 5: Performance comparison of different models on the $\text{BBBP}_{\text{reg}}^{\text{B3DB}}$ dataset with RMSE and Pearson Correlation scores. The red bars showcase the performances of our geometric graph learning-based model, GMC-MPNN. We compare against several baseline and state-of-the-art models: GCN [kipf2016semi], NF [nf], D-MPNN [heid2023chemprop], GIN [genova2017graph], GAT [velickovic2017graph], Weave [kearnes2016molecular], AttentiveFP [xiong2019pushing], CoMPT [compt], CD-MVGNN [cdmvgnn], GSL-MPP [gslmpp], and MPNN [gilmer2017neural].
Table 4: Performance of various models on the $BBBP_{reg}^{\text{B3DB}}$ dataset, sorted by ascending RMSE. PCC stands for Pearson’s correlation coefficient.
@@
[…]
We evaluate the performance of our proposed GMC-MPNN model on the regression task using the $BBBP_{\text{reg}}^{\text{B3DB}}$ dataset. GMC-MPNN, which incorporates geometry-aware WCS descriptors and max-pooling, achieves the best results among all evaluated models. It records the lowest RMSE of $0.5628\pm 0.0651$ and the highest Pearson’s correlation coefficient (PCC) of $0.6947\pm 0.0515$. The strong performance demonstrates the effectiveness of WCS features in capturing both spatial and chemical contexts critical for BBB permeability prediction.
Table 4 and Figure 5 present a comparative analysis across baseline models. Following GMC-MPNN, CoMPT [compt] emerges as the next best performer, achieving strong RMSE scores of $0.5749\pm 0.0538$. Moderate results are observed from models such as CD-MVGNN [cdmvgnn], GCN [kipf2016semi], NF [nf], D-MPNN [heid2023chemprop], and GIN [genova2017graph], record competitive RMSEs in the range of 0.62 to 0.68. Lower performance is observed from GAT [xiong2019pushing] and GSL-MPP [gslmpp], with RMSE just over 0.70. In contrast, Weave [kearnes2016molecular] and AttentiveFP [attentivefp] yield weaker outcomes, particularly AttentiveFP with a low PCC of $0.2992\pm 0.0816$. MPNN [gilmer2017neural] performs the worst, with the highest RMSE ($0.7850\pm 0.0387$) and a negative PCC ($0.0400\pm 0.1133$), indicating poor generalization.
@@ Acknowledgements
Availability of data and materials
The source code is available at the GitHub repository: https://github.com/MathIntelligence/GMC-MPNN-BBBP.